OpenStudy (anonymous):

A number is multiplied by 2,and then the product is added to 9 the result is 49 that is the number ?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

"a number"=x (x*2)+9=49

OpenStudy (anonymous):

2x+9=49 solve it for x 2x=40 x=20

OpenStudy (anonymous):

x*2=2x 2x+9=49 2x=40 x=20

OpenStudy (anonymous):

2x+9 = 49 subtract both sides by 9 u would get: 2x+9-9=49-9 2x=40 dividing both sides by 2 u would get: 2x/2=40/2 x=20